3 次元画像処理 2 単位 (選択)

Three-dimensional Image Processing  河田佳樹 · 准教授/ システム創生工学専攻光システム工学コース光情報システム工学講座

【授業目的】 3 次元画像処理の基礎理論と基本的な処理アルゴリズムを習得する ことを目的とする． 【授業概要】 3 次元画像の解析のための基礎理論と処理アルゴリズムについて解説 し，医学における画像診断支援や工業計測などへの応用例について講述する． 【授業形式】 [形態] 【キーワード】 3 次元画像処理，医用画像処理，工業計測 【先行科目】 [先行科目] 【関連科目】『バーチャルリアリティ技術』 (0.5) 【履修要件】信号処理，画像処理，プログラミング言語 (C,C++) を履修している ことが望ましいが，受講していない学生にも理解できるように配慮する． 【履修上の注意】 [注意] 【到達目標】 1. 3 次元画像処理に必要な基礎事項を理解する． 2. 医学診断や工業計測の分野で用いられている処理アルゴリズムについて理 解する． 【授業計画】 1. 3 次元画像処理概論 2. 3 次元画像処理に関する信号処理の基礎 3. 3 次元イメージング技術について 4. 3 次元画像の前処理-平滑化- 5. 3 次元画像の前処理-画像強調- 6. 3 次元画像のセグメンテーション-エッジ・領域ベースドセグメンテーション 7. 3 次元画像のセグメンテーション-変形モデルによるセグメンテーション- 8. 3 次元画像のセグメンテーション-グラフカットによるセグメンテーション- 9. 3 次元画像の幾何学的性質-画像の局所特徴量と図形のトポロジー 10. 3 次元画像の幾何学的性質-オイラー数・連結指数の計算法- 11. 3 次元画像処理のアルゴリズム-薄面化・細線化- 12. 3 次元画像処理のアルゴリズム-モルフォロジー処理- 13. 3 次元画像処理のアルゴリズム-距離変換- 14. 3 次元画像処理のアルゴリズム-微分特徴計算- 15. 3 次元画像処理のアルゴリズム-剛体・非剛体レジストレーション- 【成績評価基準】イメージング技術に関するコンピュータプロジェクトレポートで

評価する．【教科書】 3 次元ディジタル画像処理，鳥脇純一郎，昭晃堂, 2002 【参考書】

⋄ Medial Imaging Signals and Systems, J.L. Prince, J.M. Links, 2006 ⋄ image Processing, Analysis, and Machin Vision, M. Sonka, V. Hlavac, R. R. Boyle, Thomson, 2008

【WEB 頁】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=197132 【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216673 【対象学生】 [対象学生] 【連絡先】

⇒ 河田 (光棟 508, 088-656-9431, kawata@opt.tokushima-u.ac.jp) ���� 【備考】

⋄ コンピュータプロジェクトは， C または C++を用いて行う． ⋄ 授業を受ける際には， 2 時間の授業時間毎に 2 時間の予習と 2 時間の復習

をしたうえで授業を受けることが，授業の理解と単位取得のために必要である．

Three-dimensional Image Processing 2 units (selection)

Yoshiki Kawata · Associate Professor / Optical Materials and Devices, Optical Systems Engineering, Systems Innovation Engineering

Target⟩ Introduce the fundamental concepts of 3-D image processing. Outline⟩ This course is intended to introduce the fundamentals of three-dimensional

image processing which covers basic concepts as image restoration, image segmentation, registration, shape representation, and computational geometry etc. The class begins with a brief overview of the various technologies used to analyze medical and industrial images. The focus then shifts to in-depth descriptions of individual algorithms beginning with a description of the mathematical technique of image processing. The course emphasizes the design, analysis, and implementation of algorithms in the context of 3-D medical images.

Style⟩ [形態] �������⟩ 3-D image processing, medical image analysis, 工業計測 Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ “Virtual Reality”(0.5) Requirement⟩ It is desired to finish a course of digital signal processing, image

processing, and programming (C or C++). Notice⟩ [注意]

Goal⟩

1. To understand fundamentals of 3-D image processing.

2. To understand various 3-D image processing algorithms used to obtain

medical and industrial fields

Schedule⟩

1. Introduction to 3-D image processing

2. Signal Processing Fundamentals

3. 3-D imaging technology

4. 3-D image smoothing

5. 3-D image enhancement

6. Edge / region based segmentation

7. Deformable model segmentation

8. Graph cut segmentation

9. Geometrical properties of 3-D images- local feature of a connected component-

10. Geometrical properties of 3-D images- Calculation of the Euler number-

11. Surface/axis thinning algorithm 12. Morphology filter

13. Distance transformation

14. 3-D differential features

15. 3-D registration

Evaluation Criteria⟩ Computer Project Report 100%

Textbook⟩ 3 , 2002

Reference⟩

⋄ Medial Imaging Signals and Systems, J.L. Prince, J.M. Links, 2006 ⋄ image Processing, Analysis, and Machin Vision, M. Sonka, V. Hlavac, R. R. Boyle, Thomson, 2008

Webpage⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=197132 Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216673 Student⟩ [対象学生] Contact⟩

⇒ Kawata (Opt.508, +81-88-656-9431, kawata@opt.tokushima-u.ac.jp) ���� Note⟩

⋄ The computer projects should be done in C or C++.

⋄ Preparation (2hrs) and Review (2hrs) are required to take this lecture (2hrs).